OpenStudy (anonymous):

3. A surfer drives his dune buggy out into the sand dunes. He follows his compass 10 miles due north and then turns due west. If he ends up approximately 35 miles from where he started, how far west did he travel? (1 point) A) 10 miles B) 25 miles C) 33.5 miles D) 35 miles

OpenStudy (anonymous):

C

OpenStudy (anonymous):



OpenStudy (anonymous):

use Pythagorean theorem to solve for x. \[10^{2}+x ^{2}=35^{2}\]

OpenStudy (anonymous):

\[100+x ^{2}=1225\] \[100-100+x ^{2}=1225-100\] \[x ^{2}=1125\] \[\sqrt{x ^{2}}=\sqrt{1125}\] \[x=33.5\]

OpenStudy (anonymous):

Make Sense?

OpenStudy (anonymous):

yes I just forgot to do the square root.....thanks

OpenStudy (anonymous):

np